OpenStudy (sanoofy):

A trigonometric ratio is 7/5. What ratio could it be, and what angle might it be referring to? How do I even start off and do this?

OpenStudy (solomonzelman):

Well, we know that \(\sin(x)\) and \(\cos(x)\) are in \([-1,1]\) for any \(x\). (In other words, \(-1\le \sin(x)\le 1 \) and \(-1\le \cos (x)\le 1 \), for all \(x\).)

OpenStudy (solomonzelman):

There can be multiple answers .... We know that it is possible that \(\sec(x)<-1 \), \(\csc(x)<-1 \), and \(\tan(x)<-1 \).

OpenStudy (solomonzelman):

Or, it's also possible that \(\cot(x)<-1 \).

OpenStudy (solomonzelman):

So it really could be any of these 4 ...

OpenStudy (sanoofy):

Alright, thanks, I'll see what it is..